a certain monohydric aliphatic alkanol was found to contain 12g of hydrogen. determine the molecular formula
1
Expert's answer
2020-10-08T13:58:20-0400

CnH2n+1OH - formula

It's molar mass: 12n+2n+1+16+1= 14n+18

It contains 2n+2 H

The ratio is

2n+2/14n+18 = 12 / mass of alcanol

As we are not given mass of alconol, we substitute it with x

We end up with the equation:

n(x-84)+x=108

It has solutions:

n = 1, x = 96 g, methanol

n = 2, x = 92 g, ethanol

n = 3, x = 90 g, propanol

n = 4, x = 88.8 g, buthanol and so on...

So the mass of the sample is definately requirred to solve the problem :)